# Argumentative Experience: Reducing Confirmation Bias on Controversial Issues through LLM-Generated Multi-Persona Debates

### v2

@@ 3.2. Design Objectives
Figure 2. The interface of our multi-persona debate system is shown for the topic “Should human colonize space?” LLM-generated personas are shown at left, which users can freely modify, replace, or delete. Once the debate starts, thumbs up/down buttons to the top-right of each argument can be used to bookmark it for the user’s later review, recording whether or not the given argument resonated with the user’s own thinking. Bookmarked arguments are collected for the user’s review in the Summary tab.
@@ 3.3. Interface Features
Figure 2 presents the user interface of our LLM debate system. We describe the major interactive features below.
@@ 3.3.1. Controversial Topics Entry or Selection
At the start, users can either enter a topic, possibly controversial, in the search bar or choose one from the list of suggested topics A. The system then begins generating LLM-personas relevant to that topic.
@@ 3.3.2. Persona Creation
As displayed in B, each persona is represented by a title and description, along with automatically assigned colors and icons to help users easily identify and differentiate between them (which emerged as a need during early pilot testing). We set a predefined number of personas at the beginning (N=5), but users are free to generate or manually create as many personas as they want. The specific prompt instructions (described in Section 3.4.1) aim to generate diverse LLM personas that represent a range of viewpoints and social roles on the selected topic. This approach aligns with our primary design goal of comprehensive representation.
@@ 3.3.3. Interactive Debate Flow
During the design of this interface, we addressed a critical design consideration regarding the optimal level of inter-persona interaction within each debate round.
@@ 3.4.1. Persona Creation
In creating personas for our multi-persona debate system, prior studies have pointed out several limitations inherent to LLMs (Park et al., 2022, 2023b), such as inconsistent persona generation and formatting. While LLMs can generate personas that clearly express their stances on controversial topics, determining how to effectively prompt the LLM to achieve this may not always be straightforward.
One problem is that assigning specific names to personas, real or fictional, can be problematic because readers may mistakenly believe the persona represents an actual person, as evidenced by prior studies that LLMs would inadvertently reveal real names from their training data (Huang et al., 2022). To mitigate misrepresentation and potential ethical concerns, we prompt the LLM to ”generate a title for the persona, not a name”, while enabling users to override this default setting through customization. In generating persona descriptions, the LLM may include detailed background information without explicitly clarifying the persona’s stance on the topic, yet explicitly linking persona background to specific stance risks perpetuating stereotypes or creating caricatures of those backgrounds. To address this, we first prompt the LLM to “describe background (instead of their stance)” and then separately prompt it to “start the debate by briefly presenting an argument in support of its stance on the topic.”
We generate an initial set of personas to cover key perspectives on the issue. To enhance perspectival diversity, we incorporate existing personas as context when generating additional ones. Our prompt design includes specific instructions to maintain consistency across persona professions and descriptions (see Appendix A for prompt details).
@@ 3.4.2. Debate Generation
To facilitate the debate, we developed two prompts: one to initiate the debate and another to continue it based on prior debate conversations (see Appendix A for prompts details). However, it is important to note that LLMs are constrained by a maximum input length, known as the ‘context window’ (Wang et al., 2024) after which they may start to lose track of earlier portions of the input. This limitation can be problematic in our system, as personas need to reference the entire debate history to provide meaningful responses.
To overcome this challenge, inspired by Park et al. (2023b)’s work, when the debate history exceeds the LLM’s context window, we employ an extractive summarizer that compresses the transcript to fit within the required length while preserving the key points of the discussion. Specifically, we use the BERTExtractiveSummarizer (Miller, 2019), which provides a good balance between summarization performance and latency. To make the responses more concise, readable, and suitable, we also include prompt instructions “Generate at most 200 words” and “Use bullet points (in markdown format)”. This helps to reduce verbose, unstructured LLM generations and mitigate information overload.
@@ 3.4.4. Web Framework
We developed the system UI using Streamlit [^5] https://streamlit.io/, a popular machine learning web development framework that allows development with an emphasis on interactivity and ease of use. It also promotes rapid prototyping, which made it easy to iterate on the interface design based on user feedback or emerging requirements.
@@ 4. Methods
To explore how users might engage with the multi-persona debate system and the system effectiveness of reducing confirmation bias, we conducted a controlled, within-subjects study by comparing the multi-persona debate system with a baseline retrieval-based system. In this section, we describe our formal evaluation protocol, including the experimental procedure and tasks (Section 4.1), participant recruitment (Section 4.2), the selection of controversial topics (Section 4.3), and data collection and analysis.
@@ 4.1. Experimental Procedure and Tasks
As illustrated in Figure  3, the within-subject experiment started with eye calibration and task instructions. Participants were then randomly assigned to one of two study blocks, Block A or Block B. In Block A, participants utilized the baseline system to engage with the topic. We adopted ArgumentSearch [^6] https://argumentsearch.com/ (Daxenberger et al., 2020) as the baseline for our A/B experiments. ArgumentSearch is a retrieval-based search engine that identifies articles from the open web and organizes content snippets according to their stance on the user query (detailed description is provided in Appendix C).
In Block B, participants interacted with our multi-persona debate system as the experimental condition to explore and generate debate content with the given topic. Each block included a training task to first familiarize participants with the assigned system. Within each block, participants completed four randomized trials assigned with different controversial topics. Once they completed the trials in one block, they moved on to the other block and used the alternate system, ensuring they interacted both systems.
Each trial consisted of three parts: (i) Pre-task Questions (ii) Exploring the given topic with the assigned interface, and (iii) Post-task Questions. In the pre-task questions, we asked participants to explicitly rate their topic familiarity, pre-existing belief (i.e., stance) to the topic, and the confidence for their stance. For each trial, participants were presented with a controversial issue to explore the interface. After participants completing the task, we again evaluate their current stance to the topic and the confidence score, similar to the pre-task questions.
Upon completion of all eight trials across both systems, participants were required to complete a post-test questionnaire to assess their propensity for confirmation bias. Following this, a semi-structured interview was conducted to gain insight into the users’ interactions with the systems, their perception, as well as the usability and feedback. Details of the pre-, post-task questionnaires and interview questions are presented in Appendix B. The entire experimental study took approximately 1.5 to 2 hours to complete. This included 60 to 80 minutes for the experimental portion and an additional 15 to 25 minutes for the interview.
@@ 4.2. Participant recruitment
The study was conducted in a usability lab at a university, with $N=40$ participants (19 females, 18 males, and 3 non-binary identity). Particpants were recruited using convenience sampling, and were aged between 18 and 34 years (view demographic details in Appendix D). They came from a variety of academic disciplines, ranging from applied sciences, natural sciences, engineering, to social sciences, bringing diverse educational backgrounds. Participants were pre-screened for native-level English familiarity, 20/20 vision (uncorrected or corrected). Upon completion of the study, each participant was compensated.
@@ 4.3. Selection of Controversial Topics
These topics were sourced from procon.org [^7] ProCon.org_2024 covered a diverse range of categories, including health, science and technology, economics, and entertainment. These topics are formulated as yes/no proposed actions, such as ”Should abortion bans be illegal?” or ”Should space be colonized?”. The topic itself thus embeds a “pro” stance supporting the action vs. the “con” stance against it. We select topic categories to cover a range of important issues with broad societal relevance. Each topic was selected to ensure active public discussion, involvement from various groups, reliable information from credible sources, and a balanced view of arguments, without needing much specialized knowledge (view topic details in Appendix 9). In each block, one topic was selected for each category, resulting in four tasks per block and a total of eight tasks.
@@ 4.4. Data Collection and Analysis
Data collection during the experimental procedure incorporated multiple sources: pre-task questions, user interaction logs, post-task questions, system evaluation questionnaires, and interview recordings. Eye-tracking data were gathered using the Tobii TX-300 eye-tracker and processed through Tobii Pro Lab  [^8] https://www.tobii.com/products/software/behavior-research-software/tobii-pro-lab, a commercial usability and eye-tracking software that enabled raw gaze data collection, user interaction recording, and preliminary data cleaning for subsequent analysis.”
Further details about our analysis are provided below.
@@
To categorize user interaction patterns with our LLM debate system, we conducted a clustering analysis using the following behavioral data:
@@
Following the initial clustering analysis, we performed thematic coding to summarize the qualitative feedback provided. Since we compare user experience between two different interfaces, our qualitative themes focus on factors influencing participant comparisons between the A/B systems used. We present these findings in Section 5.1.
@@
A primary research objective is to evaluate the system’s effectiveness in mitigating confirmation bias in information behavior, particularly in reducing selective exposure to attitude-consistent information. Through eye-tracking data analysis, we examine engagement patterns with content that diverges from users’ prior beliefs. Drawing on confirmation bias literature (described in Section 2.2), we hypothesize that increased temporal engagement with reading opposing viewpoints may indicate reduced confirmation bias through enhanced exposure to differing perspectives.
Additionally, we incorporate two key factors in evaluating the above hypothesis based on our study design. First, we considered the impact of users’ general tendency toward confirmation bias as a psychological disposition (described in Section 2.2), to understand how this human factor may affect the system’s effectiveness. The second factor is user topic familiarity, since we intentionally selected different social topics (described in Section 4.3). We thus formulate two further sub-questions regarding how users’ eye movements might differ based on these factors.
@@
How does use of our LLM debate system vs. the baseline system impact user engagement with content? Does this effect differ based on user levels of topic familiarity and their confirmation bias tendency?
@@
Confirmation biases tendency:
@@
Eye-tracking:
Two metrics, including total fixation duration and fixation counts, were utilized to understand user engagement. Fixations were filtered to include only good-quality fixations, lasting between 100ms and 2000ms. Additionally, both fixation duration and counts were normalized relative to the length of the content.
Areas of interests:
We also label AOIs to identify participant selective exposure to attitude-consistent versus attitude-inconsistent information. For example, if the content supports the given topic and the participant also holds a supportive view, then the content would be labeled as attitude-consistent information. Fixations falling into these AOIs represent how user engage with the attitude-consistent information. Table 1 presents how we annotate attitude-consistent and attitude-inconsistent AOIs based on user pre-task belief and the stance of the content.
@@
We present the results for RQ2 in Section 5.2, focusing on user engagement of the content.
@@
We adopt a similar approach as with RQ2. Correspondingly, we formulate two sub-research questions relevant to topic familiarity and user confirmation bias tendency:
@@
How does our LLM debate system, compared to the baseline system, impact user final belief changes? Additionally, does this effect differ based on user levels of topic familiarity and their confirmation bias tendency?
User pre- and post- belief were collected separately in the pre- and post-task questions. We calculated three types of belief changes, including non-directional, bi-directional, and weighted belief changes based on user confidence, defined as follows:
Belief change:
Users are categorized into two groups of belief changes — changed vs. unchanged — based on the difference of belief measurement for each topic. User belief is measured based on a 5-point Likert scale (i.e., very supportive, somehow supportive, neutral, somehow opposed, and very opposed), collected in the pre- and post-task questions.
Bi-directional belief change:
Users are categorized into three groups of belief changes — less supportive, more supportive, and unchanged — showing the directional belief change for each topic, as calculated based on the previous 5-point Likert scale.
Weighted belief change:
We calculate the stance shift weighted by the average confidence level before and after using each system (Equation 1) [^9] In prior behavioral studies by Norman and Price (2015), relative confidence of people’s subjective responses were evaluated using average measures. Building on this, we evaluate participants’ relative confidence in their beliefs about the topics by calculating the average confidence level before and after the intervention.. This score reflects both the direction and strength of the belief change. User belief confidence is also measured by a 5-point Likert scale in the pre- and post-task questions.
(1) $\frac{(\textnormal{Belief}_{\textnormal{after}}-\textnormal{Belief}_{\textnormal{before}})(\textnormal{Confidence}_{\textnormal{after}}+\textnormal{Confidence}_{\textnormal{before}})}{2}$

### v3

@@ 3.2. Design Objectives
Figure 2. The interface of our multi-persona debate system is shown for the topic “Should humans colonize space?” LLM personas are shown on the left, which users can freely modify, replace, or delete. Once the debate starts, thumbs up/down buttons to the top-right of each argument can be used to bookmark it for the user’s later review, recording whether or not the given argument resonated with the user’s own thinking. Bookmarked arguments are collected for the user’s review in the Summary tab.
[…]
The interface of the multi-persona debate system
@@ 3.3. Interface Features
Figure 2 presents the user interface of the LLM multi-persona debate system. We describe the major interactive features below.
@@ 3.3.1. Controversial Topics Entry or Selection
At the start, users can either enter a topic, possibly controversial, in the search bar or choose one from the list of suggested topics A. The system then begins generating LLM personas relevant to that topic.
@@ 3.3.2. Persona Creation
As displayed in B, each persona is represented by a title and description, along with automatically assigned colors and icons to help users easily identify and differentiate between them (which emerged as a need during early pilot testing). We set a predefined number of personas at the beginning (N=5). These personas are created automatically by an LLM based on the system prompt: “Given the topic, [TOPIC], create a roundtable debate of different personas to show key perspectives of the topic.” Each persona is associated with its profession and description. This prompt instruction (described in Section 3.4.1) generates diverse LLM personas. Additionally, to align with our primary design goal of comprehensive representation, users are free to generate or manually create as many personas as they want or edit the existing personas in the interface.
@@ 3.3.3. Interactive Debate Flow
During the design of this interface, we addressed the design consideration regarding the optimal level of inter-persona interaction within each debate round.
@@ 3.4.1. Persona Creation
To satisfy our first design objective (i.e., broad representation), we identified two key challenges from prior research and our pilot implementations.
First, prior work generated personas by constraining them based on different characteristics, such as interest (e.g., hobbies and jobs) and personal traits (e.g., kind, bully) (described in Section 2.1). But, we found that this method introduces a tradeoff for our study. Overly detailed specifications (e.g., exhaustive lists of traits) risk producing personas that might limit LLMs’ ability to generate nuanced and diverse perspectives. Conversely, oversimplified personas with minimum specifications may inadvertently amplify LLMs’ inherent biases, particularly those predisposed to favor dominant cultural or demographic traits (Salminen et al., 2024; Beck et al., 2023). This suggests that failing to specify appropriate traits can lead to an imbalanced spectrum of viewpoints, thereby not achieving our goals.
Another problem is that while assigning specific names, real or fictional, to personas can enhance authenticity and elicit more genuine user responses, it risks misleading readers into perceiving personas as representations of real individuals. This concern is amplified by evidence from prior studies showing that LLMs may inadvertently leak real names from their training data (Huang et al., 2022).
To address diversity limitations, we prompt the LLM to use multiple perspectives to reflect a topic and then generate personas. This ensures that even for topics like “Should abortion be banned?” that might favor certain sides, the resulting personas can capture a balanced spectrum of viewpoints, representing both supporting and opposing arguments. Additionally, we incorporate existing personas as context when users initiate the generation of a new persona. The prompt includes instructions to maintain consistency across persona professions and descriptions.
To mitigate misrepresentation and potential ethical concerns, we prompt the LLM to “generate a title for the persona, not a name” while enabling users to override this default setting through customization. In generating persona descriptions, the LLM may include detailed background information without explicitly clarifying the persona’s stance on the topic, yet explicitly linking persona background to stance risks perpetuating stereotypes or creating caricatures of those backgrounds. To address this, we first prompt the LLM to “describe background (instead of their stance)” and then separately prompt it to “start the debate by briefly presenting an argument in support of its stance on the topic.”
@@ 3.4.2. Debate Generation
To facilitate the debate, we developed two prompts: one to initiate the debate and another to continue it based on prior debate conversations. However, it is important to note that LLMs are constrained by a maximum input length, known as the ‘context window’ (Wang et al., 2024), after which they may start to lose track of earlier portions of the input. This limitation can be problematic in our system, as personas need to reference the entire debate history to provide meaningful responses.
To overcome this challenge, inspired by Park et al. (2023b)’s work, when the debate history exceeds the LLM context window, we employ an extractive summarizer that compresses the transcript to fit within the required length while preserving the key points of the discussion. Specifically, we use the BERTExtractiveSummarizer (Miller, 2019), which provides a good balance between summarization performance and latency. To make the responses more concise, readable, and suitable, we also include prompt instructions “Generate at most 200 words” and “Use bullet points (in markdown format)”. This helps to reduce verbose, unstructured LLM generations and mitigate information overload.
@@ 3.4.4. Web Framework
We developed the system UI using Streamlit [^5] https://streamlit.io/, a popular machine learning web development framework that allows development with an emphasis on interactivity and ease of use. It also promotes rapid prototyping, which makes it easy to iterate on the interface design based on user feedback or emerging requirements.
@@ 4. Methods
To explore how users might engage with the multi-persona debate system and the system’s effectiveness in reducing confirmation bias, we conducted a controlled, within-subjects study by comparing the multi-persona debate system with a baseline retrieval-based system. In this section, we describe our formal evaluation protocol, including the experimental procedure and tasks (Section 4.1), participant recruitment (Section 4.2), the selection of controversial topics (Section 4.3), and data collection and analysis.
@@ 4.1. Experimental Procedure and Tasks
[…]
Flowchart of the experimental procedure
The study followed a within-subjects design to reduce individual variability. As each participant interacted with both the baseline system and the multi-persona debate system, this approach allows for direct comparison while controlling for differences between individuals. To minimize potential order effects, both the task topics and the sequence of system exposure were randomized across participants.
As illustrated in Figure  3, the experiment started with eye calibration and task instructions. Participants were then randomly assigned to one of two study blocks, Block A or Block B. In Block A, participants utilized the baseline system to engage with the topic. We adopted ArgumentSearch [^6] https://argumentsearch.com/ (Daxenberger et al., 2020) as the baseline for our A/B experiments. ArgumentSearch is a retrieval-based search engine that identifies articles from the open web and organizes content snippets according to their stance on the user query (detailed description is provided in Appendix C).
The design of ArgumentSearch explicitly organizes arguments by stance and exposes users to multiple viewpoints from diverse web sources (Daxenberger et al., 2020). Compared to traditional search engines, which have proven to inadvertently reinforce confirmation bias (discussed in Section 2.2), the structured presentation of ArgumentSearch makes it a stronger baseline to compare with our multi-persona debate system, enabling a rigorous evaluation of our system effectiveness in mitigating confirmation bias.
In Block B, participants interacted with our multi-persona debate system as the experimental condition to explore and generate debate content with the given topic. Each block included a training task to first familiarize participants with the assigned system. Within each block, participants completed four randomized trials assigned with different controversial topics. Once they completed the trials in one block, they moved on to the other block and used the alternate system, ensuring they interacted with both systems.
Each trial consisted of three parts: (i) Pre-task Questions (ii) Exploring the given topic with the assigned interface, and (iii) Post-task Questions. In the pre-task questions, we asked participants to explicitly rate their topic familiarity, pre-existing belief (i.e., stance) of the topic, and confidence in their stance. For each trial, participants were presented with a controversial topic to explore the interface. After participants complete the task, we again evaluate their current stance on the topic and the confidence score, similar to the pre-task questions.
Upon completion of all eight trials across both systems, participants were required to complete a post-test questionnaire to assess their propensity for confirmation bias. Following this, a semi-structured interview was conducted to gain insight into the users’ interactions with the systems, their perception, and the usability and feedback. Details of the pre- and post-task questionnaires and interview questions are presented in Appendix B. The entire experimental study took approximately 1.5 to 2 hours to complete. This included 60 to 80 minutes for the experimental portion and an additional 15 to 25 minutes for the interview.
@@ 4.2. Participant Recruitment
The study was conducted in a usability lab at a university, with $N=40$ participants (19 females, 18 males, and 3 non-binary identities). Participants were recruited using convenience sampling and were aged between 18 and 34 years (view demographic details in Appendix D). They came from a variety of academic disciplines, ranging from applied sciences, natural sciences, and engineering to social sciences, bringing diverse educational backgrounds. Participants were pre-screened for native-level English familiarity and 20/20 vision (uncorrected or corrected). The study was approved by the Institutional Review Board (IRB), and each participant was compensated with 30 USD upon completion.
@@ 4.3. Selection of Controversial Topics
These topics were sourced from procon.org [^7] ProCon.org_2024 and covered a diverse range of categories, including health, science and technology, economics, and entertainment. These topics are formulated as yes/no proposed actions, such as ”Should abortion bans be illegal?” or ”Should space be colonized?”. The topic itself thus embeds a “pro” stance supporting the action vs. the “con” stance against it.
We chose these topics as they reflect polarized political discourses currently. Participants might hold strong prior beliefs about them, making confirmation bias more likely to influence how they process information. These topics provide a relevant context for evaluating whether the multi-persona debate system can promote open-minded engagement and effectively mitigate biased attention to attitude-consistent information. Each topic was selected to ensure active public discussion, involvement from various groups, reliable information from credible sources, and a balanced view of arguments without needing much specialized knowledge (view topic details in Appendix 9). In each block, one topic was selected for each category, resulting in four tasks per block and a total of eight tasks.
To ensure the quality and stance alignment of the generated arguments, each persona was prompted with structured instructions to reflect its assigned viewpoint on the topic. We also conducted post-hoc annotations to verify that each persona consistently expressed its intended stance. Personas identified as holding a neutral stance were excluded from the data analysis focused on quantifying patterns of user interaction with attitude-consistent versus attitude-inconsistent arguments. In the multi-persona debate system design, users are allowed to freely select personas representing diverse stances. To address potential imbalances in stance exposure, we normalized the collected behavioral data, particularly the eye-tracking measures, by accounting for both the number of pro and con personas and the length of the generated debate content. This normalization reduces exposure effects and enables more accurate comparisons of user engagement across different viewpoints.
@@ 4.4. Data Collection and Analysis
Data collection during the experimental procedure incorporated multiple sources: pre-task questions, user interaction logs and eye movements, post-task questions, system evaluation questionnaires, and interview recordings. Eye-tracking approach offers objective insights into users’ visual attention and engagement with system content, particularly in assessing how confirmation bias manifests during interaction. Eye-tracking data were gathered using the Tobii TX-300 eye-tracker and processed through Tobii Pro Lab  [^8] https://www.tobii.com/products/software/behavior-research-software/tobii-pro-lab, a commercial usability and eye-tracking software that enabled raw gaze data collection, user interaction recording, and preliminary data cleaning for subsequent analysis.
Eye-tracking data and behavior logs were filtered to exclude segments involving task scenario reading and pre/post-task questionnaires, retaining only the periods of active interface interaction. For analysis, we removed low-quality and non-reading eye-tracking data by discarding fixations shorter than 100 ms or longer than 1500 ms. After preprocessing, statistical assumptions were assessed for all analyses reported in the paper. Additionally, normality tests (mainly Shapiro-Wilk and Q-Q plots) were performed on the dependent variables, including eye-tracking measures and user belief change data, followed by the use of parametric or non-parametric tests. Further analysis details are provided below.
@@
To answer this question, we employed both quantitative and qualitative data analysis. We first analyzed the interaction patterns that participants developed when using the system and correlated them with qualitative data. A clustering analysis was used to uncover interaction patterns based on the following behavioral data :
@@
Following the initial clustering analysis, we performed thematic coding to summarize the qualitative feedback provided. Since we compare user experience between two different interfaces, our qualitative themes focus on factors influencing participant comparisons between the A/B systems used. Two rounds of qualitative coding were conducted, each performed by different researchers. In the first round, two researchers employed bottom-up thematic open coding without referencing the quantitative data, ensuring an unbiased data exploration. In the second round, a third researcher performed thematic axial coding based on the clustered groups.
The above analysis resulted in two main findings to answer the first question: (1) an integrated result combining quantitative and qualitative insights that illustrate participants’ behavioral patterns, and (2) a set of qualitative themes that capture overarching user behaviors when engaging with the tool. We present them in Section 5.1.
@@
A primary research objective is to evaluate the system effectiveness in mitigating confirmation bias, particularly in reducing selective exposure to attitude-consistent information. Drawing on cognitive literature studying confirmation bias user engagement — specifically, heightened attention to content aligning with users’ prior beliefs — serves as a cognitive manifestation of confirmation bias (Betsch et al., 2001; Jonas et al., 2008) (the Motivational aspect of confirmation bias, described in Section 2.2). To investigate this, we leverage eye-tracking data to analyze engagement patterns with attitude-inconsistent information (i.e., viewpoints diverging from users’ existing stances). We hypothesize that prolonged dwell times on such content may signal reduced confirmation bias, as extended exposure to opposing viewpoints encourages cognitive reevaluation.
Additionally, we incorporate two key independent variables (IV) in evaluating the above hypothesis based on our study design. First, we considered the impact of users’ general tendency toward confirmation bias as a psychological disposition (the Cognitive aspect of confirmation bias, described in Section 2.2), investigating how this human factor affects the system effectiveness. The second variable is user topic familiarity since we intentionally selected different social topics (described in Section 4.3).
Corresponding eye-tracking data is our dependent variable (DV). We thus formulate two further sub-questions regarding how users’ visual attention might differ based on these factors.
@@
How does the use of LLM multi-persona debate system vs. the baseline system impact user engagement with content? Does this effect differ based on user levels of topic familiarity and their confirmation bias tendency?
@@
IV: Confirmation biases tendency:
@@
DV: Eye-tracking:
Two metrics were utilized to understand user engagement, including total fixation duration and fixation counts. Fixations were filtered to include only good-quality fixations, lasting between 100ms and 2000ms. Additionally, both fixation duration and counts were normalized relative to the length of the content.
DV: Areas of interests:
We label AOIs to identify participant selective exposure to attitude-consistent versus attitude-inconsistent information. For example, if the content supports the given topic and the participant also holds a supportive view, then the content would be labeled as attitude-consistent information. Fixations falling into these AOIs represent how users engage with the attitude-consistent information. Table 1 presents how we annotate attitude-consistent and attitude-inconsistent AOIs based on user pre-task belief and the stance of the content.
@@
We present the results for RQ2 in Section 5.2
@@
We adopt a similar approach as with RQ2. Note that a reduced confirmation bias, as manifested in user engagement of attitude-inconsistent information, might not finally result in user belief changes. To validate this relationship, we formulate two sub-research questions relevant to the prior IV: topic familiarity and user confirmation bias tendency:
@@
How does the LLM multi-persona debate, compared to the baseline system, impact user final belief changes? Additionally, does this effect differ based on user levels of topic familiarity and their confirmation bias tendency?
To address both questions, we measure user belief changes (i.e., the difference between user pre- and post-belief) as another DV. User pre- and post-belief were collected separately in the pre- and post-task questions. We calculated three types of belief changes, including non-directional, directional, and weighted belief changes based on user confidence, defined as follows:
DV: Belief change:
Users are categorized into two groups of belief changes — changed vs. unchanged (Equation 1) — based on the difference of belief measurement for each topic. User belief is measured based on a 5-point Likert scale (i.e., very supportive, somehow supportive, neutral, somehow opposed, and very opposed), collected in the pre- and post-task questions.
DV: Directional belief change:
Since the prior belief change does not indicate whether users become more supportive or opposed to the topic, we use a directional measure to differentiate both. We categorize the directional belief change into — less supportive, more supportive, and unchanged (Equation 2).
DV: Weighted belief change:
Furthermore, we calculate the belief change weighted by the average confidence level before and after using each system (Equation 3) [^9] In prior behavioral studies by Norman and Price (2015), relative confidence of people’s subjective responses were evaluated using average measures. Building on this, we evaluate participants’ relative confidence in their beliefs about the topics by calculating the average confidence level before and after the intervention.. This score reflects both the direction and strength of changes. User belief confidence is also measured by a 5-point Likert scale in the pre- and post-task questions.
(1) $Belief\ change=\text{sgn}(|Belief_{after}-Belief_{before}|)$
(2) $Belief\ change_{directional}=\text{sgn}(Belief_{after}-Belief_{before})$
(3) $Belief\ change_{weighted}=\frac{\text{(}Belief_{after}-Belief_{before})(Confidence_{after}+Confidence_{before})}{2}$